# Explore-Go: Leveraging Exploration for Generalisation in Deep Reinforcement Learning

### v1

@@ 2.1 Contextual Markov decision process
A contextual MDP (CMDP, Hallak et al., 2015) is a specific type of MDP where the state space $S=S^{\prime}\times C$ can in principle be factored into an underlying state space $S^{\prime}$ and a context space $C$. In other words, a state $s=(s^{\prime},c)\in S$ in a CMDP can be split into an underlying state $s^{\prime}\in S^{\prime}$ and a context $c\in C$. The context $c$ behaves differently than the underlying state $s^{\prime}$ in that it is sampled at the start of an episode (as part of the distribution $p_{0}$) and remains fixed until the episode ends. The context $c$ can be thought of as the task an agent has to solve and from here on out we will refer to context and task interchangeably.
@@ 6.1 Generalisation in CMDPs
So far we have mentioned some approaches that increase the number and variability of the training tasks. Other approaches instead try to explicitly bridge the gap between the training and testing tasks. For example, some use inductive biases to encourage learning generalisable functions (Zambaldi et al., 2018, 2019; Kansky et al., 2017; Wang et al., 2021; Tang et al., 2020; Tang & Ha, 2021). Others use regularisation techniques from supervised learning to boost generalisation performance (Cobbe et al., 2019; Tishby & Zaslavsky, 2015; Igl et al., 2019; Lu et al., 2020; Eysenbach et al., 2021). We mention only a selection of approaches here, for a more comprehensive overview we refer to the survey by Kirk et al. (2023).
@@ 6.2 Exploration in CMDPs
Closest to our work is Weltevrede et al. (2023), Jiang et al. (2023) and Suau et al. (2023). Weltevrede et al. (2023) introduce the concepts of reachable generalisation but don’t provide intuition on what to do for unreachable generalisation, nor do they propose a novel, scalable approach for maximising generalisation. Jiang et al. (2023) don’t make a distinction between reachable and unreachable generalisation and provide intuition which we argue mainly applies to reachable generalisation (see Appendix A). Moreover, their novel approach only works for off-policy algorithms, whereas ours could be applied to both off-policy and on-policy methods. The concurrent work in Suau et al. (2023) introduces the notion of policy confounding in out-of-trajectory generalisation. The issue of policy confounding is complementary to our intuition for unreachable generalisation. However, it is unclear how out-of-trajectory generalisation equates to reachable or unreachable generalisation. Moreover, they do not propose a novel, scalable approach to solve the issue.
@@ 7 Conclusion
One of the limitations of our approach arises if the pure exploration agent is very inefficient. If this is the case, Explore-Go tends to waste interactions with the environment that are not used by the main agent to improve its performance. For future work, we propose to improve the pure exploration agent and investigate ways to increase the starting state distribution of the agent with as little additional environment interactions as possible, for example, by leveraging off-policy algorithms like DQN (Mnih et al., 2015) or TD3 (Fujimoto et al., 2018).

### v3

@@ 2.1 Contextual Markov decision process
A contextual MDP (CMDP, Hallak et al., 2015) is a specific type of MDP where the state space $S=S^{\prime}\times C$ can in principle be factored into an underlying state space $S^{\prime}$ and a context space $C$. For a state $s=(s^{\prime},c)\in S$, the context $c$ behaves differently than the underlying state $s^{\prime}$ in that it is sampled at the start of an episode (as part of the distribution $p_{0}$) and remains fixed until the episode ends. The context $c$ can be thought of as the task an agent has to solve and from here on out we will refer to context and task interchangeably.
@@ 6.1 Generalisation in CMDPs
So far we have mentioned some approaches that increase the number and variability of the training tasks. Other approaches instead try to explicitly bridge the gap between the training and testing tasks. For example, some use inductive biases to encourage learning generalisable functions (Zambaldi et al., 2018, 2019; Kansky et al., 2017; Wang et al., 2021; Tang et al., 2020; Tang & Ha, 2021).
Others use regularisation techniques from supervised learning to boost generalisation performance (Cobbe et al., 2019; Tishby & Zaslavsky, 2015; Igl et al., 2019; Lu et al., 2020; Eysenbach et al., 2021). We mention only a selection of approaches here, for a more comprehensive overview we refer to the survey by Kirk et al. (2023).
@@ 6.2 Exploration in CMDPs
In Zisselman et al. (2023), the authors leverage exploration at test time to move the agent towards states where it can confidently solve the task, thereby increasing test time performance. Our work differs in that we leverage exploration during training time in order to increase the number of states from which the agent can confidently solve the test tasks. Closest to our work is Weltevrede et al. (2023), Jiang et al. (2023), Zhu et al. (2020) and Suau et al. (2023). Weltevrede et al. (2023) introduce the concepts of reachable generalisation but don’t provide intuition on what to do for unreachable generalisation, nor do they propose a novel, scalable approach for maximising generalisation. Jiang et al. (2023) don’t make a distinction between reachable and unreachable generalisation and provide intuition which we argue mainly applies to reachable generalisation (see Appendix A). Moreover, their novel approach only works for off-policy algorithms, whereas ours could be applied to both off-policy and on-policy methods. In Zhu et al. (2020), the authors learn a reset controller that increases the diversity of the agent’s start states. However, they only argue (and empirically show) that this benefits reachable generalisation. The concurrent work in Suau et al. (2023) introduces the notion of policy confounding in out-of-trajectory generalisation. The issue of policy confounding is complementary to our intuition for unreachable generalisation. However, it is unclear how out-of-trajectory generalisation equates to reachable or unreachable generalisation. Moreover, they do not propose a novel, scalable approach to solve the issue.
@@ 7 Conclusion
One of the limitations of our approach arises if the pure exploration agent is very inefficient. If this is the case, Explore-Go tends to waste interactions with the environment that are not used by the main agent to improve its performance. For future work, we propose to improve the pure exploration agent and investigate ways to increase the starting state distribution of the agent with as little additional environment interactions as possible, for example, by leveraging an off-policy algorithms like DQN (Mnih et al., 2015) or a model-based approach like PEG (Hu et al., 2023).
Acknowledgements
This work was partially funded by the Dutch Research Council (NWO) project Reliable Out-of-Distribution Generalization in Deep Reinforcement Learning with project number OCENW.M.21.234.